OpenStudy (evansanchez):

Which two points are on the graph of y = x - 3? (1, -2), (-1, 3) (1, 2), (0, 3) (0, -3), (4, 1) (3, 0), (0, 3) Is it C?

OpenStudy (skullpatrol):

please go back and look at this answer

jimthompson5910 (jim_thompson5910):

You are correct again. -------------------- Check plug in (x,y) = (0,-3) y = x-3 -3 = 0-3 -3 = -3 ... works So (0,-3) is on the line plug in (x,y) = (4,1) y = x-3 1 = 4-3 1= 1 ... works So (4,1) is on the line